You are buying a house and will be taking out a traditional 30-year home mortgage loan for $210,000.

(a) Plan A, with its APR=5.75%, below. Show work. (3 pts)

Monthly Payment = $


Total Paid = $
1
Expert's answer
2020-05-06T16:50:56-0400


  1. find an effective bid

"EAR=(1+\frac{0.0575}{12})^{12}-1=0.05903987"

2.find the monthly rate

"(1+r)^{12}+1=1+EAR"

r=1.05903987^1/12-1=0.00479167

2.find the monthly payment

"A=\frac{210 000\times0.00479167}{1-1.00479167^{-360}}"


A=1 225.50

Monthly Payment = $1 225.50

"1225.50\times360=441180"

Total Paid = $441 180